### v1

@@ 4.2 The correct formula for quantifying uncertainty
where $Q_{j}=\sum_{i\neq j}v_{i}^{\prime}A_{ij}$ is mean-zero and independent of $(e_{j},v_{j})$. The above is the so-called Hoeffding decomposition (Hoeffding, 1948); the two sums are the first two U-statistics, which are uncorrelated with each other by construction. Thus, the variance decomposes as:
@@ 5.2 Technical details
Let $\xi_{i}$ be independent $T_{i}\times 1$ random vectors with $\mathbb{E}[\xi_{i}]=0$, $\mathbb{V}(\xi_{i})=I_{T_{i}}$, and $\mathbb{E}|\tau^{\prime}\xi_{i}|^{4}\leq C|\tau|^{4}$ for all $T_{i}\times 1$ vectors $\tau$. Let $\Gamma$ be an $n\times n$ block matrix with blocks $\Gamma_{ij}$ of size $T_{i}\times T_{j}$, symmetric across the diagonal ($\Gamma_{ij}=\Gamma_{ji}^{\prime}$), and with $\Gamma_{ii}=0$ for all $i$. Assume:
@@
The assumptions above ensure two linear CLTs: one for sums of $e_{i}$ and one for sums of $v_{i}\otimes e_{i}$. Condition (ii) guarantees that $(\Sigma_{v,i}^{-1/2}v_{i})\otimes(\Sigma_{e,i}^{-1/2}e_{i})$ has eigenvalues bounded away from zero, preventing near-degeneracy. Condition (iii) ensures that correlations between $e_{i}$ and $v_{i}\otimes e_{i}$ remain bounded away from one. The moment conditions in (iv) rule out heavy-tailed behavior that could dominate the covariance-based approximation.
@@
Part (i) guarantees that the normalization in the CLT is no slower than $1/\sqrt{n}$. Part (ii) ensures cluster negligibility for the linear CLT, Part (iii) (b) does so for the quadratic CLT, while Part (iii) (a) assumes that quadratic part is negligible in comparison to the linear one.
@@
Suppose the data is indexed by $\ell=\ell(i,t)$ and that a standard weak exogeneity condition holds: $\mathbb{E}[e_{it}\!\mid\!x_{is}:s\leq t]=0$. Let $\mathcal{X}_{0}=\{\ell(i,1)\}$ denote the set of first-period observations for each cluster. For these $\ell$, we have $\mathcal{E}_{\ell\tilde{\ell}}=1$ for all $\tilde{\ell}$, so no observations are excluded from the projection, and thus $A^{*}_{\ell\ell}=M_{\ell\ell}$. If $\min_{\ell\in\mathcal{X}_{0}}M_{\ell\ell}>c>0$ and average cluster size is bounded, then Assumption 4(i) holds. The condition $M_{\ell\ell}>c$ is standard in the many regressors/instruments literature (e.g., Cattaneo et al. (2018)). $\Box$
@@
For linear CLTs, the condition $\max_{i}T_{i}^{2}/n\to 0$ suffices (Hansen and Lee, 2019; Sasaki and Wang, 2022), while the quadratic CLT (Assumption 4(iii)(b)) requires the stricter condition $\max_{i}T_{i}^{4}/n\to 0$.
In high-dimensional settings with many regressors, an analog of the second part of Assumption 4(ii)—e.g., $\frac{\|M\lambda\|_{\infty}^{2}}{n}\to 0$—is often imposed as a high-level assumption because direct verification is difficult; see, for example, Assumption 3 in Cattaneo et al. (2018). Lemma 7 offers low-level primitive conditions that imply this assumption. The bound $\|\lambda\|_{\infty}<\infty$ is mild since $\lambda_{\ell}=\mathbb{E}[x_{\ell}]$, and for a fixed effects projection matrix, we have $\|M\|_{\infty}=2$.
@@ 6.1 Jackknife variance estimator
The jackknife variance estimator, introduced by Efron and Stein (1981) for symmetric statistics based on independent and identically distributed data, has also been advocated as a useful tool in settings involving non-symmetric statistics and non-identically distributed data (see, e.g., Bousquet et al., 2004; MacKinnon et al., 2023). In our setting, we define the statistic of interest as $\mathcal{Z}=x^{\prime}A^{*}U=x^{\prime}A^{*}e=f(D_{1},\dots,D_{N})$, where $D_{i}=(x_{i},U_{i})$ denotes the observed data for cluster $i$. Let $\mathcal{Z}_{(i)}=f(D_{1},\dots,D_{i-1},0,D_{i+1},\dots,D_{N})$ denote the statistic computed when cluster $i$ is removed (i.e., replaced by zero). The jackknife estimator of the variance is then given by
@@
Efron and Stein (1981) showed that the jackknife variance estimator tends to be conservative for symmetric statistics derived from independent and identically distributed data. While the jackknife correctly captures the variance of first-order $U$-statistics, it double-counts second-order terms. In our setting, $\mathcal{Z}$ is non-symmetric and based on independent but non-identically distributed data. The jackknife is conservative for two main reasons: it gives double weight to the quadratic term (as noted in Efron and Stein (1981)), and $\mathcal{Z}_{(i)}$ is improperly centered since $\mathbb{E}[U_{i}-e_{i}]\neq 0$. The jackknife variance estimator is unbiased when $A^{*}_{ij}=0$ for all $i\neq j$, in which case the removal of a single cluster does not affect the residualization of others. Consequently, $\mathcal{Z}_{(i)}$ is properly centered, and no cross-cluster covariances arise. In general, the degree of conservativeness is small when $\sum_{j=1}^{N}\sum_{i\neq j}\|A_{ij}^{*}\|_{F}^{2}$ is much smaller than $\sum_{j=1}^{N}\|A_{jj}^{*}\|_{F}^{2}$.
@@ 6.2 Tests and confidence sets
Our estimator, $\hat{\beta}^{A^{*}}=\frac{x^{\prime}A^{*}y}{x^{\prime}A^{*}x}=\frac{z^{\prime}y}{z^{\prime}x}$, belongs to the class of just-identified internal IV estimators. Similar to the IV estimator for dynamic panels proposed by Arellano and Bond (1991), our approach may be subject to weak identification concerns (see Bun and Windmeijer (2010)). To construct inference procedures that are robust to weak identification, we propose using the Anderson–Rubin (AR) test. The AR test delivers valid inference under both strong and weak identification.
@@ 7 Empirical Application
Policy interventions with spatial effects present a compelling setting for applying our method. In such contexts, researchers often distinguish between the direct effect of treatment—on the treated units themselves—and the total effect, which includes indirect impacts arising from spillovers. A prominent example is provided by Egger et al. (2022) (hereafter EHMNW), who study a large-scale fiscal intervention in rural Kenya. In their experiment, one-time cash transfers in total equivalent to approximately 15 percent of local GDP were randomly assigned to 653 villages. We apply our method to this setting to estimate the direct effect of the intervention on treated villages.
@@
This setting fits naturally within our design-based framework, which we illustrate through estimation under two specifications. In both cases, we define the exogeneity matrix $\mathcal{E}$ using cutoffs based on pairwise distances between villages. We impose the exclusion restriction $E[v_{\tilde{\ell}}({y}_{\ell}-\beta x_{\ell})]=0$ if and only if the distance between villages $\tilde{\ell}$ and $\ell$ is at least $R$ kilometers. In practice, researchers should choose a distance cutoff below which exogeneity may be questionable. Here, we illustrate the behavior of the $\hat{\beta}^{A^{*}}$ estimator using a range of cutoffs from 0 to 3 kilometers. In the data, the average pairwise distance between villages within a sub-location is $2.2$ kilometers. On average, each village has $1.5$, $4.6$, and $6.7$ neighbors within the sub-location and within 1, 2, and 3 kilometers [^3] Based on GPS coordinates. We dropped nine observations for which we do not have GPS coordinates., respectively. Our estimates under different choices of $R$ are shown in Figure 1.
@@
For specification (a), let ${y}_{\ell}$ denote the average consumption of households in village $\ell$, and let $x_{\ell}$ denote the binary treatment status of the village. We assume the correctly specified treatment equation: $x_{\ell}=p_{i(\ell)}+v_{\ell},$ where $p_{i}$ is a cluster (sub-location) fixed effect capturing treatment propensity. We include sub-location fixed effects because the realized treatment share can differ significantly from the ex ante assignment probability, especially in smaller sub-locations.
@@
Specification (b), reported in Figure 1, Panel (b), makes use of the full variation in treatment intensity. Here, we define ${y}_{\ell}$ as the average consumption of eligible households in village $\ell$, and let $x_{\ell}$ denote the total transfer allocated to the village. The assumed treatment assignment equation in this case is $x_{\ell}=W_{\ell}^{\prime}\delta+v_{\ell},$ where $W_{\ell}$ includes the number of eligible households in village $\ell$ along with sub-location fixed effects. The parameter of interest, $\beta$, again represents the village-level direct treatment effect. This parameter is somewhat comparable to the total effect on eligible households reported in EHMNW, which was estimated using an instrumental variables strategy and reported in Table 1 (row 1, column 2) of the paper. Estimates under different exogeneity assumptions are presented along with both jackknife and cluster-robust standard errors. In this specification, the $A^{*}$ matrix is no longer block-diagonal, so the two variance estimators do not coincide.

### v2

@@ 4.2 The correct formula for quantifying uncertainty
where $Q_{j}=\sum_{i\neq j}v_{i}^{\prime}A_{ij}$ is mean-zero and independent of $(e_{j},v_{j})$. The above is the so-called Hoeffding decomposition (hoeffding1948class); the two sums are the first two U-statistics, which are uncorrelated with each other by construction. Thus, the variance decomposes as:
@@ 5.2 Technical details
Let $\xi_{i}$ be independent $T_{i}\times 1$ random vectors with $\mathbb{E}[\xi_{i}]=0$, $\mathbb{V}(\xi_{i})=I_{T_{i}}$, and $\mathbb{E}|\tau^{\prime}\xi_{i}|^{4}\leq C\|\tau\|^{4}$ for all $T_{i}\times 1$ vectors $\tau$. Let $\Gamma$ be an $n\times n$ block matrix with blocks $\Gamma_{ij}$ of size $T_{i}\times T_{j}$, symmetric across the diagonal ($\Gamma_{ij}=\Gamma_{ji}^{\prime}$), and with $\Gamma_{ii}=0$ for all $i$. Assume:
@@
The assumptions above ensure two linear CLTs: one for sums of $e_{i}$ and one for sums of $v_{i}\otimes e_{i}$. Condition (ii) guarantees that $(\Sigma_{v,i}^{-1/2}v_{i})\otimes(\Sigma_{e,i}^{-1/2}e_{i})$ has eigenvalues bounded away from zero, preventing near-degeneracy. Condition (iii) ensures that correlations between $e_{i}$ and $v_{i}\otimes e_{i}$ remain bounded away from one. Condition (iv) rules out heavy-tailed behavior and strong higher-order dependence that could dominate the covariance-based approximation.
@@
Part (i) guarantees that the normalization in the CLT is no slower than $1/\sqrt{n}$. Part (ii) ensures cluster negligibility for the linear CLT, Part (iii) (b) does so for the quadratic CLT, while Part (iii) (a) assumes that the quadratic part is negligible in comparison to the linear one.
@@
Suppose the data is indexed by $\ell=\ell(i,t)$ and that a standard weak exogeneity condition holds: $\mathbb{E}[e_{it}\!\mid\!x_{is}:s\leq t]=0$. Let $\mathcal{X}_{0}=\{\ell(i,1)\}$ denote the set of first-period observations for each cluster. For these $\ell$, we have $\mathcal{E}_{\ell\tilde{\ell}}=1$ for all $\tilde{\ell}$, so no observations are excluded from the projection, and thus $A^{*}_{\ell\ell}=M_{\ell\ell}$. If $\min_{\ell\in\mathcal{X}_{0}}M_{\ell\ell}>c>0$ and average cluster size is bounded, then Assumption 4(i) holds. The condition $M_{\ell\ell}>c$ is standard in the many regressors/instruments literature (e.g., cattaneo2018inference). ∎
@@
For linear CLTs, the condition $\max_{i}T_{i}^{2}/n\to 0$ suffices (hansen2019asymptotic; sasaki2022non), while the quadratic CLT (Assumption 4(iii)(b)) requires the stricter condition $\max_{i}T_{i}^{4}/n\to 0$.
In high-dimensional settings with many regressors, an analog of the second part of Assumption 4(ii)—e.g., $\frac{\|M\lambda\|_{\infty}^{2}}{n}\to 0$—is often imposed as a high-level assumption because direct verification is difficult; see, for example, Assumption 3 in cattaneo2018inference. Lemma 7 offers low-level primitive conditions that imply this assumption. The bound $\|\lambda\|_{\infty}<\infty$ is mild since $\lambda_{\ell}=\mathbb{E}[x_{\ell}]$, and for a fixed effects projection matrix, we have $\|M\|_{\infty}=2$.
@@ 6.1 Jackknife variance estimator
The jackknife variance estimator, introduced by efron1981jackknife for symmetric statistics based on independent and identically distributed data, has also been advocated as a useful tool in settings involving non-symmetric statistics and non-identically distributed data (see, e.g., bousquet2004concentration; mackinnon2023fast). In our setting, we define the statistic of interest as $\mathcal{Z}=x^{\prime}A^{*}U=x^{\prime}A^{*}e=f(D_{1},\dots,D_{N})$, where $D_{i}=(x_{i},U_{i})$ denotes the observed data for cluster $i$. Let $\mathcal{Z}_{(i)}=f(D_{1},\dots,D_{i-1},0,D_{i+1},\dots,D_{N})$ denote the statistic computed when cluster $i$ is removed (i.e., replaced by zero). The jackknife estimator of the variance is then given by
@@
efron1981jackknife showed that the jackknife variance estimator tends to be conservative for symmetric statistics derived from independent and identically distributed data. While the jackknife correctly captures the variance of first-order $U$-statistics, it double-counts second-order terms. In our setting, $\mathcal{Z}$ is non-symmetric and based on independent but non-identically distributed data. The jackknife is conservative for two main reasons: it gives double weight to the quadratic term (as noted in efron1981jackknife), and $\mathcal{Z}_{(i)}$ is improperly centered since $\mathbb{E}[U_{i}-e_{i}]\neq 0$. The jackknife variance estimator is unbiased when $A^{*}_{ij}=0$ for all $i\neq j$, in which case the removal of a single cluster does not affect the residualization of others. Consequently, $\mathcal{Z}_{(i)}$ is properly centered, and no cross-cluster covariances arise. In general, the degree of conservativeness is small when $\sum_{j=1}^{N}\sum_{i\neq j}\|A_{ij}^{*}\|_{F}^{2}$ is much smaller than $\sum_{j=1}^{N}\|A_{jj}^{*}\|_{F}^{2}$.
@@ 6.2 Tests and confidence sets
Our estimator, $\hat{\beta}^{A^{*}}=\frac{x^{\prime}A^{*}y}{x^{\prime}A^{*}x}=\frac{z^{\prime}y}{z^{\prime}x}$, belongs to the class of just-identified internal IV estimators. Similar to the IV estimator for dynamic panels proposed by arellano1991some, our approach may be subject to weak identification concerns (see bun2010weak). To construct inference procedures that are robust to weak identification, we propose using the Anderson–Rubin (AR) test. The AR test delivers valid inference under both strong and weak identification.
@@ 7 Empirical Application
Policy interventions with spatial effects present a compelling setting for applying our method. In such contexts, researchers often distinguish between the direct effect of treatment—on the treated units themselves—and the total effect, which includes indirect impacts arising from spillovers. A prominent example is provided by egger2022general (hereafter EHMNW), who study a large-scale fiscal intervention in rural Kenya. In their experiment, one-time cash transfers in total equivalent to approximately 15 percent of local GDP were randomly assigned to 653 villages. We apply our method to this setting to estimate the direct effect of the intervention on treated villages.
@@
This setting fits naturally within our design-based framework, which we illustrate through estimation under two specifications. In both cases, we define the exogeneity matrix $\mathcal{E}$ using cutoffs based on pairwise distances between villages. We assume the exclusion restriction $E[v_{\tilde{\ell}}({y}_{\ell}-\beta x_{\ell})]=0$ holds if and only if $\ell=\tilde{\ell}$ or the distance between villages $\tilde{\ell}$ and $\ell$ is at least $R$ kilometers. In practice, researchers should choose a distance cutoff below which exogeneity may be questionable. Here, we illustrate the behavior of the $\hat{\beta}^{A^{*}}$ estimator using a range of cutoffs from 0 to 3 kilometers. In the data, the average pairwise distance between villages within a sub-location is $2.2$ kilometers. On average, each village has $1.5$, $4.6$, and $6.7$ neighbors within the sub-location and within 1, 2, and 3 kilometers [^3] Based on GPS coordinates. We dropped nine observations for which we do not have GPS coordinates., respectively. Our estimates under different choices of $R$ are shown in Figure 2.
@@
For specification (a), let ${y}_{\ell}$ denote the average consumption of households in village $\ell$, and let $x_{\ell}$ denote the binary treatment status of the village. We assume the correctly specified treatment equation: $x_{\ell}=p_{i(\ell)}+v_{\ell},$ where $p_{i}$ is a cluster (sub-location) fixed effect capturing treatment propensity. We include sub-location fixed effects to flexibly absorb the realized treatment share, which may differ substantially from the ex ante assignment probability. In practice, the realized propensity can depend on cluster size or arise from re-randomization or re-balancing procedures implemented during the experiment.
@@
Specification (b), reported in Figure 2, Panel (b), makes use of the full variation in treatment intensity. Here, we define ${y}_{\ell}$ as the average consumption of eligible households in village $\ell$, and let $x_{\ell}$ denote the total transfer allocated to the village. The assumed treatment assignment equation in this case is $x_{\ell}=W_{\ell}^{\prime}\delta+v_{\ell},$ where $W_{\ell}$ includes the number of eligible households in village $\ell$ along with sub-location fixed effects. The parameter of interest, $\beta$, again represents the village-level direct treatment effect. This parameter is somewhat comparable to the total effect on eligible households reported in EHMNW, which was estimated using an instrumental variables strategy and reported in Table 1 (row 1, column 2) of the paper. Estimates under different exogeneity assumptions are presented along with both jackknife and cluster-robust standard errors. In this specification, the controls $W_{\ell}$ include a covariate that varies within clusters and is not absorbed by the sub-location fixed effects. As a result, matrix $M$ is no longer block-diagonal, and neither is $A^{*}$, so the two variance estimators do not coincide.
Discussion.
Both panels illustrate our main takeaway: the point estimates—and especially their precision—are highly sensitive to the researcher’s maintained exogeneity assumptions. Relaxing these assumptions, for example by allowing spillovers to extend within 3 km rather than 2 km, leads to substantially wider confidence intervals. This reflects a reduction in the effective sample size captured by the trace of the $A^{*}$ matrix under less restrictive exogeneity assumptions.